How much water must be added to 75 ml of a 75% alcohol solution to obtain 45% alcohol solution?
Answer by rvquartz(19)

let x be the amount of water to add
The final amount of liquid will be (75 + x)
the key thing to realize is that the amount of actual pure acid before adding water is THE SAME as the amount of actual pure acid after adding water - its just that in one case you have less liquid but a stronger solution and in the other case you have more liquid but a weaker solution

since the amount of pure acid is the same we can say that
0.75(75) = 0.45(75 + x)
56.25 = 33.75 + 0.45x
if a = b, then b = a, so
33.75 + 0.45x = 56.25
0.45x = 22.5
x = 22.5/0.45
x = 50 ml